The annual income from an apartment complex is $21439. The annual expense is estimated to be $3181. The apartment complex could be sold for $129301 at the end of 10 years. If your MARR is 10%, how much should you pay for the apartment complex if you were to buy it now?
Answer #1

ANSWER:

Annual income = $21,439

Annual expense = $3,181

Annual benefit = Annual income - Annual expense = $21,439 - $3,181 = $18,258

Salvage value = $129,301

n = 10 years

i = 10%

Present value = Annual benefit(p/a,i,n) + salvage value(p/f,i,n)

Present value = 18,258(p/a,10%,10) + 129,301(p/f,10%,10)

Present value = 18,258 * 6.145 + 129,301 * 0.3855

Present value = 112,195.41 + 49,845.535

Present value = 162,040.95

So, right now he should pay $162,040.95 for the house.